OpenStudy (anonymous):

Please oh please help!!! Admission to the zoo is $6 for children and $9 for adults. On a certain day, 1480 people go to the zoo and $10,500 is collected. How many children and how many adults went to the zoo? ______ Children ______ Adults

OpenStudy (anonymous):

6x+9y = 10500 x+y = 1480 Using elimination I get: 6x + 9y = 10500 -6x-6y = -8880 Add the 2 equations and I get 3y = 1620 Divide both sides by 3 and I get y = 540 Substitute that back in to 6x+9y = 10500. I get that 6x+9*540 = 10500 9 * 540 = 4860 Subtract 4860 from both sides and I get 6x = 5640 Divide both sides by 6 and I get x = 940 So there must have be 940 children and 540 adults

OpenStudy (anonymous):

Can you help with another question please?